Calculate the gauge pressures (in N/m2) inside 2.14 cm radius bubbles of the following liquids.

water

alcohol

soapy water
Answer #1

For a bubble with two surfaces, the pressure difference between the inside and the outside is
pi - pf = 4 / r
Where is the surface tension and r is the radius of the bubble.
1) For water
the surface tension = 0.073 N/m
Difference between the pressure is called the gauge pressure
The gauge pressure = 4 / r
The gauge pressure = 4 x 0.073 / 2.14 x 10-2 m
The gauge pressure =  13.6  N/m2
2) For alcohol
Surface tension = 0.022
The gauge pressure = 4 / r
The gauge pressure = 4 x 0.022 / 2.14 x 10-2 m
The gauge pressure =  4.11 N/m2
3) For soapy water,
surface tension = 0.025
The gauge pressure = 4 / r
The gauge pressure = 4 x 0.025 / 2.14 x 10-2 m
The gauge pressure =  4.67 N/m2